OpenStudy (anonymous):

Could someone help double-check 5 shell integration problems please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks, used washer method on 3 and 4

OpenStudy (anonymous):

@israa88 you seen anything wrong?

OpenStudy (anonymous):

give me a min plz

OpenStudy (anonymous):

I am not 10 percent sure but it seems good to me

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem :)

OpenStudy (amistre64):

first one seems wrong

OpenStudy (amistre64):

wait, you doing shells on 1 ... let me look at it again

OpenStudy (amistre64):

int 2pir h r = x, h = x^(1/2) 2pi int[0,4] x^(3/2) dx setups fine

OpenStudy (amistre64):

2nd one, tell me about your limits .... you did a u sub but i dont see you changing back to x, but you are using the limits of x

OpenStudy (amistre64):

the interval is defined as 0 < x < sqrt(pi) ; let u = x^2 0^2 < x^2 < sqrt(pi)^2 0 < u < pi your solution is not calculated in the correct domain

OpenStudy (anonymous):

i didn't explicitly change to x in a step, but did so in plugging in the bounds

OpenStudy (anonymous):

I did think that the area enclosed could be wrong, in that there really is no left bounds by the description of the problem and the graph could continue indefinitely to the left

OpenStudy (amistre64):

redo #2, its not organized correctly and it has a few flaws

OpenStudy (anonymous):

is the graph correct?

OpenStudy (amistre64):

the graph is fine :) the rest is a muck

OpenStudy (anonymous):

ok, 3,4,5, look ok?

OpenStudy (anonymous):

ahh, I see that the bounds for 1 should be based on u after subbing

OpenStudy (amistre64):

\[\int_{x=0}^{x=\sqrt{\pi}}2\pi~r~h\] \[2\pi\int_{x=0}^{x=\sqrt{\pi}}~x~sin(x^2)~dx\] \[2\pi\int_{x=0}^{x=\sqrt{\pi}}~\frac12sin(u=\color{red}{x^2})~du=\color{red}{2xdx}\] \[2\pi\int_{x=0}^{x=\sqrt{\pi}}~\frac12sin(u)~du\] \[\pi\int_{\color{red}{x=0}}^{\color{red}{x=\sqrt{\pi}}}~sin(u)~du\]

OpenStudy (amistre64):

u(x) = x^2 u(0) = 0^2 = 0 u(sqrt pi) = (sqrt pi)^2 = pi

OpenStudy (amistre64):

it wasnt as amuck as i originally thought :) but that was due to you leaving some stuff out of it. define u and du for your substitutions, they are not clearly defined

OpenStudy (anonymous):

i see

OpenStudy (amistre64):

you have a sideway picture for 3, is that acceptable? just wondering

OpenStudy (amistre64):

why would you change your method on #3? did you have to do washers?

OpenStudy (amistre64):

is it user choice? i mean its a fine approach just curious

OpenStudy (amistre64):

x^(1/3) integrates to : x^(1/3 + 3/3) ------------- 1/3 + 3/3 does this match your integration?

OpenStudy (amistre64):

4 is off, you are spose to revolve about x=4, not y=4

OpenStudy (anonymous):

ahh, thanks for that catch

OpenStudy (anonymous):

user choice on number 3

OpenStudy (amistre64):

\[\int_{0}^{2}2pi ~rh\] \[\int_{0}^{2}2pi ~(4-x)(2\sqrt2~x^{1/2}-x^2)~dx\]

OpenStudy (amistre64):

why is r = (4-x) in this case?

OpenStudy (anonymous):

i'm sure I messed that up, and need to recalculate most likely, I used the washer approach, and revolved it around y = 4, need to change it to x = 4

OpenStudy (amistre64):

yeah, and use shells ...

OpenStudy (anonymous):

guess I'm not seeing on #3 why it's turned on its side, the graph seems correct

OpenStudy (amistre64):